### v1

@@ I Introduction
Modeling activities are often more complex than an atomic model transformation (MT) and rely on sequences of MTs. Pertinent examples can be found in model synchronization [marussy2018incremental], model refactoring [mokaddem2018recommending], and rule-based design-space exploration [abdeen2014multi]. Typically, there might be more than one MT sequence that can successfully transform the source model into the target state, and choosing the most appropriate (cost-effective, efficient, safe) one manually is not tractable. This raises the need for automated methods for developing complex MTs, in which MTs are chained in sequences.
@@
Reinforcement learning (RL) is a machine learning (ML) paradigm, in which an agent explores the state space, and through trial and error, learns beneficial action sequences [sutton2018reinforcement]. This ability positions RL as a potential automation method for developing complex MTs. The benefits or RL have been recognized in MDE before, e.g., in model repair [barriga2019personalized], in-place MTs [eisenberg2021towards], and model-driven service migration [dehghani2022facilitating]. However, RL methods perform poorly in the early phases of learning due to suboptimal decisions before converging to better policies [sutton2018reinforcement]. That is, RL agents may require substantial training time and resources before learning an optimal sequence of MTs. This shallow learning curve limits the applicability of RL, particularly in cases where early mistakes can be costly, e.g., synchronizing models in model-based design of safety-critical systems. While the performance of RL agents improves over time, this improvement often comes too late to be useful.
@@
Human guidance is an apt technique to mitigate problems, such as the shallow learning curve of RL agents [najar2021reinforcement]. In such a setup, human advice acts as a heuristic to hint the otherwise autonomous agent towards desirable goals, thereby improving its performance. Involving human guidance in automated MDE methods has been explored before, e.g., driving RL-based MTs for model repair by personalized preferences [barriga2022parmorel], and driving search-based methods for MT-based design space exploration [hegedus2015model]. A common theme of existing techniques is the assumption that firm advice can be produced by the human in a timely and cost-effective fashion. This is a strong assumption that cannot be always guaranteed, especially when the human is asked to provide guidance about complex MTs (e.g., identifying consistent intermediate states in a long refactoring, or known locally optimal configurations in design-space exploration). Human advice, though often uncertain, can still be of high value in driving RL agents [dagenais2024towards]. This raises the need for incorporating human advice into RL-driven MTs through sound modeling and quantification of uncertainty.
@@
In this paper, we develop an approach and technical framework to infer complex MTs through RL, guided by potentially uncertain human advice. The sound formalization and assessment of uncertainty allows for producing useful human advice early in the MT learning process. Depending on the specific situation and engineering problem, early uncertain advice might be of higher value than late certain advice. We evaluate our approach in synthetic and real human guidance scenarios. Our results show that human guidance, even if uncertain, substantially improves RL performance, and allows for the efficient inference of complex MTs.
@@ II-A Reinforcement learning
In RL, an agent learns optimal control of an environment by taking sequential actions to explore the environment and update its strategy based on feedback in the form of rewards [sutton2018reinforcement]. This can be formalized as a Markov decision process [puterman1990markov] $\langle S,A,\pi(a|s),r_{t}(s,a)\rangle$. $S$ denotes the set of observable states the environment can produce and $A$ is the set of actions the agent can take. $r_{t}(s,a)\in\mathbb{R}$ is the reward the environment produces when the agent chooses action $a\in A$ while observing state $s\in S$, at time step $t$. At time step $t$, the agent observes a state $s_{t}$ and chooses an action $a_{t}$. In response, at time step $t+1$, the environment produces a reward $r_{t+1}$ and a new state $s_{t+1}$. The agent maintains a policy $\pi(a|s)$, which gives the conditional probability of choosing action $a$ in state $s$. The agent’s goal is to learn and act according to an optimal policy $\pi_{*}$ that maximizes the sum of future rewards.
RL has seen growing use in support of MDE, e.g., in MT engineering [eisenberg2021towards], model repair [barriga2022parmorel], fault injection testing [moradi2020exploring], and inference of simulators for digital twins [david2022devs]. In this work, we use RL to learn complex MTs, i.e., long sequences of MTs that are infeasible to reason about for human modelers.
@@ II-B Human guidance in RL
[…]
@@ II-C Uncertainty and Subjective Logic
Using human guidance to bias the RL agent’s policy requires a formal approach to appropriately capture human opinions. In particular, the epistemic uncertainty of human opinions must be modeled appropriately to make guidance meaningful. Subjective Logic (SL) [j_osang2016subjective] is a formalism that defines an opinion, a construct composed of belief and an orthogonal component of uncertainty. A binomial opinion is defined as $\omega_{x}=(b_{x},d_{x},u_{x},a_{x})$. This opinion is about the truth of a boolean predicate $x$. Here $b_{x}$ is belief in $x$, $d_{x}$ disbelief in $x$, $u_{x}$ vacuity of evidence of $x$, and $a_{x}$ prior probability of $x$. The parameters must adhere to the constraints $b_{x},d_{x},u_{x},a_{x}\in[0,1]$ and $b_{x}+d_{x}+u_{x}=1$. A binomial opinion is transformed into a probability $P(x)=b_{x}+a_{x}u_{x}$. Likewise, probability $p(x)$ translates to binomial opinion $\omega=(p,1-p,0,p)$. SL defines fusion operators that allow for multiple opinions to be fused into one opinion.
The utility of SL in MDE has been demonstrated, e.g., in modeling with uncertain types [burgueno2018expressing, bagheri2009belief, troya2021uncertainty] and prioritizing inconsistency resolution steps [jongeling2023uncertainty]. In this work, we use SL to model the uncertainty of guidance for RL in MT development.
@@ II-D Related Work
Closest to our work are approaches that use RL for MTs, particularly with human in the loop for guidance or intervention. The PARMOREL framework [barriga2022parmorel, iovino2020model] uses RL to learn user-preferred sequences of model transformations to fix invalid models. While this approach features human input for choosing an appropriate reward function, it does not accommodate uncertainty of human guidance.
@@
eisenberg2021towards [eisenberg2021towards, eisenberg2024single] use RL for learning in-place MTs. In this approach, the RL agent learns a sequence of MTs that optimizes objectives in the final model. The authors demonstrate that RL is a feasible alternative to traditional search heuristics, e.g., genetic algorithms. Our work improves over these approaches by including human guidance in RL.
@@
basciani2018tool develop the CITRIC tool to recommend developers optimal MT sequences between a source and target metamodel. The approach ranks MT sequences through shortest path algorithms by two relevant criteria: metamodel coverage and information loss. Our approach is similar in aims but differs in technological solution, including the role of the human, who is an active advisor in our approach.
[…]
III Illustrative example
To illustrate our approach, we rely on the following running example of a self-driving vehicle through a frozen lake (Fig. 1). The example is analogous to Open AI’s Gym’s Frozen Lake environment [^1] https://gymnasium.farama.org/environments/toy_text/frozen_lake/, but it is fully modeled in EMF and its operational semantics are defined by MTs, akin to the Pacman case, a frequently used example in MDE [heckel2006graph, syriani2019domain-specific, eisenberg2021towards].
(a) Metamodel
@@ III Illustrative example
Operational semantics are given by MTs. As shown in Fig. 2, moving to the tile to right is achieved by executing the moveRightRule MT rule, detailed in Listing 1. To illustrate our approach, we use the Viatra framework [bergmann2015viatra].
@@
Through trial-and-error, the agent gradually learns which actions are beneficial in a specific situation. Eventually, it will develop a policy that allows for navigating from Start to Goal without falling into Holes. By that, the policy, encodes the MT sequence of interest, which can be materialized by selecting the actions (MTs) with the highest probability in each state.
The complexity of the problem prevents the agent from learning at high pace. Typically, the agent struggles at the beginning of the learning process [sutton2018reinforcement], especially when obstacles hinder exploration (e.g., the Hole in [1, 1], and the Holes on the right edge forming a narrow passage through [2, 1]).
@@
The advisor might be not completely certain in the advice. For example, advisors might be more certain about the vicinity of their location (situational expertise), or might be experts in recognizing Holes but not Goals (domain-specific expertise).
@@ IV-B Domain-specific (meta)modeling
The framework allows for the execution MTs as RL programs in an MDE process, with minimal development effort.
@@
Related works have provided clear mapping between MTs and RL [eisenberg2021towards, barriga2019personalized]. Here, we only recall that graph morphisms or MT left-hand sides (LHS) correspond to states in RL programs; and model manipulations correspond to actions in RL. Reward is automatically calculated (e.g., number of violations in model repair [barriga2019personalized]), or assumes a reward structure (e.g., quality of solutions in model exploration [abdeen2014multi]).
@@ IV-D Operational semantics
If the agent is situated in a terminating state at this point, the lifecycle continues with Terminating the episode, which includes updating the policy an Starting a new episode by resetting the model (e.g., placing the agent in the initial state). This internal MT is shown in Listing 3 and Listing 4.
@@
If the agent is situated in a non-terminating state, the execution continues with recording the reward for the successful step and matching LHSs in the new state.
@@ V Guiding RL agents by opinions
In this section, we elaborate on Step 4. Our previous work [dagenais2024opinion, dagenais2024towards] presents the general method in great detail. Here, we focus on adopting its essential elements required for this paper. We refer to Fig. 5 in the remainder of this section.
@@ V-A Advisor input
Advice is provided through an appropriate domain-specific language (DSL) [schmidt2006model] and ideally, in a Syntax-directed editor. To demonstrate principles and conduct our experiments, we developed a DSL and an Xtext-based editor for the running example. Such editors can be developed or generated for the specific case at hand. The DSL to provide advice about the grid world of the running example is shown in Listing 5. An example set of advice expressed in the DSL about the running example (Fig. 1) is shown in Listing 6.
@@ V-B Compiling advice into opinion
Next, in step 4.2, advice is transformed into opinions by mapping to the parameters of subjective logic [josang2016subjective]—briefly introduced in Sec. II-C. This step is completely automated and is achieved in three sub-steps as follows.
@@ V-B1 Calibrating base rate
In the running example, there are four actions the agent can choose in any state, thus $|A|=4$ and the base rate $a_{i}=0.25$. (Note that if the agent chooses an action that would lead it out of the grid world, it stays in place)
@@ V-B2 Calibrating uncertainty
Next, we calibrate uncertainty $u_{i}$. While uncertainty can be set manually by the advisor, it can prove difficult to quantify, and thus we recommend deriving uncertainty using an automated method. For example, uncertainty may be derived from an appropriate distance metric, such that uncertainty increases with distance. To calibrate uncertainty using a distance measure it is essential to choose (i) an appropriate distance metric and (ii) a discount function that successively discounts uncertainty as distance increases.
In general the discount function is given as $\gamma:(u_{max},\delta)\mapsto u_{i}\in(0,1)$, where $\delta$ is a distance measured by a distance measure $\Delta$, and $u_{max}$ is the upper bound of uncertainty. From the identity of $u_{i}=1-(b_{i}+d_{i})$ (Sec. II), it follows that $u_{max}=1-(b_{i}+d_{i})$. In the most simple case, a linear discount function can be chosen to calibrate uncertainty as follows.
@@
In problems where uncertainty should be scaled only to a subset of the problem space, we use a threshold $10<\tau\in\Delta<$ 1 to accelerate the discounting of certainty and reach $u_{max}$ in $\tau$. Then, Equation 2 is adopted as follows.
@@ V-B3 Computing disbelief (d) and belief (b) to form an opinion
$\begin{split}\omega_{i}&=\left(\frac{j-1}{n-1}\times(1-u_{i}),\frac{n-j}{n-1}\times(1-u_{i}),u_{i},\frac{1}{|A|}\right)\\ &=\left(\frac{5-1}{5-1}\times(1-0.5),\frac{5-5}{5-1}\times(1-0.5),0.5,\frac{1}{|4|}\right)\\ &=\left(0.5,0,0.5,0.25\right).\\ \end{split}$
@@ V-C1 Transforming Agent’s Policy into certainty domain
To fuse opinions described in the certainty domain, with a policy described in the probability domain, we choose to transform the agent’s policy to the certainty domain (step 4.3). Our choice is motivated by the well-defined and sound fusion operators that are readily available in the toolbox of subjective logic [josang2016subjective]. This translation is achieved as follows: $f_{P\rightarrow\Omega}(\pi):S\times A\rightarrow\Omega~{}|~{}P.$ As explained in Sec. II-C, probability $p$ translates to opinion $\omega:p\mapsto(p,1-p,0,p).$
@@ V-C2 Policy shaping by opinion fusion
[…]
@@ V-D Transforming opinion-infused policies from the certainty domain to the probability domain
The agent’s policy now reflects the advice and the agent may begin the exploration process.

### v2

@@ I Introduction
Modeling activities are often more complex than an atomic model transformation (MT) and rely on sequences of MTs. Pertinent examples can be found in model synchronization [1], model refactoring [2], and rule-based design-space exploration [3]. Typically, there might be more than one MT sequence that can successfully transform the source model into the target state, and choosing the most appropriate (cost-effective, efficient, safe) one manually is not tractable. This raises the need for automated methods for developing complex MTs, in which MTs are chained in sequences.
@@
Reinforcement learning (RL) is a machine learning (ML) paradigm, in which an agent explores the state space, and through trial and error, learns beneficial action sequences [4]. This ability positions RL as a potential automation method for developing complex MTs. The benefits of RL have been recognized in MDE before, e.g., in model repair [5], in-place MTs [6], and model-driven service migration [7]. However, RL methods perform poorly in the early phases of learning due to suboptimal decisions before converging to better policies [4]. That is, RL agents may require substantial training time and resources before learning an optimal sequence of MTs. This shallow learning curve limits the applicability of RL, particularly in cases where early mistakes can be costly, e.g., synchronizing models in model-based design of safety-critical systems. While the performance of RL agents improves over time, this improvement often comes too late to be useful.
@@
Human guidance is an apt technique to mitigate problems, such as the shallow learning curve of RL agents [8]. In such a setup, human advice acts as a heuristic to hint the otherwise autonomous agent towards desirable goals, thereby improving its performance. Involving human guidance in automated MDE methods has been explored before, e.g., driving RL-based MTs for model repair by personalized preferences [9], and driving search-based methods for MT-based design space exploration [10]. A common theme of existing techniques is the assumption that firm advice can be produced by the human in a timely and cost-effective fashion. This is a strong assumption that cannot always be guaranteed, especially when the human is asked to provide guidance about complex MTs (e.g., identifying consistent intermediate states in a long refactoring, or known locally optimal configurations in design-space exploration). Human advice, though often uncertain, can still be of high value in driving RL agents [11]. This raises the need for incorporating human advice into RL-driven MTs through sound modeling and quantification of uncertainty.
@@
In this paper, we develop an approach and technical framework to infer complex MTs through RL, guided by potentially uncertain human advice. The sound formalization and assessment of uncertainty allows for producing useful human advice early in the MT learning process. Depending on the specific situation and engineering problem, early uncertain advice might be of higher value than late certain advice. We evaluate our approach in synthetic and real human guidance scenarios. Our results indicate that human guidance, even if uncertain, substantially improves RL performance, and allows for the efficient inference of complex MTs. [^1] A replication package with detailed replication instructions, results, and analyses is publicly available at https://doi.org/10.5281/zenodo.16321499.
@@ II-A Reinforcement learning
[…]
RL has seen growing use in support of MDE, e.g., in MT engineering [6], model repair [9], fault injection testing [13], and inference of simulators for digital twins [14]. In this work, we use RL to learn complex MTs, i.e., long sequences of MTs that are infeasible to reason about for human modelers.
@@ II-B Human guidance in RL
While RL algorithms can operate autonomously, evidence suggests that human guidance is beneficial to learning dynamics. Guidance is a method in which human input is used to bias the agent’s exploration strategy. Techniques differ by when and how advice is provided, as described in the taxonomy of [8]. For example, aligned with our approach, [15] emphasize the role of early advice.
Strategies to incorporate guidance into RL are called policy shaping methods, which are defined by which part of the RL process they are incorporated in. In value shaping, human advice acts as a function of action preferences [16]. In reward shaping, advice is translated into numerical rewards given to the agent [17]. In decision biasing, human advice is used to directly affect the policy output [18]. Finally, in policy shaping, human advice is directly incorporated into the agent’s policy to bias the exploration strategy. For example, [19] estimate the human’s feedback policy and incorporate this information into the agent’s policy.
@@ II-C Uncertainty and Subjective Logic
Using human guidance to bias the RL agent’s policy requires a formal approach to appropriately capture human opinions. In particular, the epistemic uncertainty of human opinions must be modeled appropriately to make guidance meaningful. Subjective Logic (SL) [20] is a formalism that defines an opinion, a construct composed of belief and an orthogonal component of uncertainty. A binomial opinion is defined as $\omega_{x}=(b_{x},d_{x},u_{x},a_{x})$. This opinion is about the truth of a Boolean predicate $x$. Here $b_{x}$ is belief in $x$, $d_{x}$ disbelief in $x$, $u_{x}$ vacuity of evidence of $x$, and $a_{x}$ prior probability of $x$. The parameters must adhere to the constraints $b_{x},d_{x},u_{x},a_{x}\in[0,1]$ and $b_{x}+d_{x}+u_{x}=1$. A binomial opinion is transformed into a probability $P(x)=b_{x}+a_{x}u_{x}$. Likewise, probability $p(x)$ translates to binomial opinion $\omega=(p,1-p,0,p)$. SL defines fusion operators that allow for multiple opinions to be fused into one opinion.
The utility of SL in MDE has been demonstrated, e.g., in modeling with uncertain types [21, 22, 23] and prioritizing inconsistency resolution steps [24]. In this work, we use SL to model the uncertainty of guidance for RL in MT development.
@@ II-D Related Work
Closest to our work are approaches that use RL for MTs, particularly with human in the loop for guidance or intervention. The PARMOREL framework [9, 25] uses RL to learn user-preferred sequences of model transformations to fix invalid models. While this approach features human input for choosing an appropriate reward function, it does not accommodate the uncertainty of human guidance.
[6] [6, 26] use RL for learning in-place MTs. In this approach, the RL agent learns a sequence of MTs that optimizes objectives in the final model. The authors demonstrate that RL is a feasible alternative to traditional search heuristics, e.g., genetic algorithms. Our work improves over these approaches by including human guidance in RL.
[27] introduce the CITRIC tool to recommend developers optimal MT sequences between a source and target metamodel. The approach ranks MT sequences through shortest path algorithms by two relevant criteria: metamodel coverage and information loss. Our approach is similar in aims but differs in technological solution, including the role of the human, who is an active advisor in our approach.
Finally, some adjacent techniques that automate MT development, but not by RL, include the following. [28] use multi-objective optimization to explore MT sequences using meta-heuristic search algorithms. [3] use a genetic algorithm for multi-objective optimization in rule-based design-space exploration. [2] learn refactoring rules from examples using genetic programming.
III Illustrative example
To illustrate our approach, we rely on the following running example of a self-driving vehicle through a frozen lake (Fig. 1). The example is analogous to Open AI’s Gym’s Frozen Lake environment [^2] https://gymnasium.farama.org/environments/toy_text/frozen_lake/, but it is fully modeled in EMF and its operational semantics are defined by MTs, akin to the Pacman case, a frequently used example in MDE [29, 30, 6].
@@ III Illustrative example
[trim=7cm 6cm 6cm 6.5cm,clip,width=]figures/lake-class-diagram.pdf
(a) Metamodel
@@
[width=]figures/default-annotated.pdf
@@
Operational semantics are given by MTs. As shown in Fig. 2, moving to the tile to the right is achieved by executing the moveRightRule MT rule, detailed in Listing 1. To illustrate our approach, we use the Viatra framework [31].
[…]
[trim=0.75cm 0.2cm 0.75cm 0.1cm,clip,width=]figures/mt-lake-horizontal.pdf
@@
Through trial-and-error, the agent gradually learns which actions are beneficial in a specific situation. Eventually, it will develop a policy that allows for navigating from Start to Goal without falling into Holes. By that, the policy encodes the MT sequence of interest, which can be materialized by selecting the actions (MTs) with the highest probability in each state.
The complexity of the problem prevents the agent from learning at a high pace. Typically, the agent struggles at the beginning of the learning process [4], especially when obstacles hinder exploration (e.g., the Hole in [1, 1], and the Holes around [2, 1], forming a narrow passage).
@@
The advisor might not be completely certain in the advice. For example, advisors might be more certain about the vicinity of their location (situational expertise), or might be experts in recognizing Holes but not Goals (domain-specific expertise).
@@ IV-B Domain-specific (meta)modeling
The framework allows for executing MTs as RL programs in an MDE process, with minimal development effort.
@@
Related works have provided a clear mapping between MTs and RL [6, 5]. Here, we only recall that graph morphisms or MT left-hand sides (LHS) correspond to states in RL programs and model manipulations correspond to actions in RL. Reward is automatically calculated (e.g., number of violations in model repair [5]), or assumes a reward structure (e.g., quality of solutions in model exploration [3]).
@@ IV-D Operational semantics
If the agent is situated in a terminating state at this point, the lifecycle continues by Terminating the episode, which includes updating the policy and Starting a new episode by resetting the model (e.g., placing the agent in the initial state). This internal MT is shown in Listing 3 and Listing 4.
@@
If the agent is situated in a non-terminating state, the execution continues by recording the reward for the successful step and matching LHSs in the new state.
@@ V Guiding RL agents by opinions
In this section, we elaborate on Step 4. Our previous work [32, 11] presents the general method in great detail. Here, we focus on adopting its essential elements required for this paper. We refer to Fig. 5 in the remainder of this section.
@@ V-A Advisor input
Advice is provided through an appropriate domain-specific language (DSL) [33] and ideally, in a Syntax-directed editor. To demonstrate principles and conduct our experiments, we developed a DSL and an Xtext-based editor for the running example. Such editors can be developed or generated for the specific case at hand. The DSL to provide advice about the grid world of the running example is shown in Listing 5. An example set of advice expressed in the DSL about the running example (Fig. 1) is shown in Listing 6.
@@ V-B Compiling advice into opinion
Next, in step 4.2, advice is transformed into opinions by mapping to the parameters of subjective logic [34]—briefly introduced in Sec. II-C. This step is completely automated and is achieved in three sub-steps as follows.
@@ V-B1 Calibrating base rate
In the running example, there are four actions the agent can choose in any state, thus $|A|=4$ and the base rate $a_{i}=0.25$. (Note that if the agent chooses an action that would lead it out of the grid world, it stays in place.)
@@ V-B2 Calibrating uncertainty
Next, we calibrate uncertainty $u_{i}$. While uncertainty can be set manually by the advisor, it can prove difficult to quantify, and thus, we recommend deriving uncertainty using an automated method. For example, uncertainty may be derived from an appropriate distance metric, such that uncertainty increases with distance. To calibrate uncertainty using a distance measure, it is essential to choose (i) an appropriate distance metric and (ii) a discount function that successively discounts uncertainty as distance increases.
In general the discount function is given as $\gamma:(u_{max},\delta)\mapsto u_{i}\in(0,1)$, where $\delta$ is a distance measured by a distance measure $\Delta$, and $u_{max}$ is the upper bound of uncertainty. From the identity of $u_{i}=1-(b_{i}+d_{i})$ (Sec. II), it follows that $u_{max}=1-(b_{i}+d_{i})$. In the simplest case, a linear discount function can be chosen to calibrate uncertainty as follows.
@@
In problems where uncertainty should be scaled only to a subset of the problem space, we use a threshold $0<\tau\in\Delta<1$ to accelerate the discounting of certainty and reach $u_{max}$ in $\tau$. Then, Equation 2 is adopted as follows.
@@ V-B3 Computing disbelief (d) and belief (b) to form an opinion
[…]
@@ V-C1 Transforming Agent’s Policy into certainty domain
To fuse opinions described in the certainty domain, with a policy described in the probability domain, we choose to transform the agent’s policy to the certainty domain (step 4.3). Our choice is motivated by the well-defined and sound fusion operators that are readily available in the toolbox of subjective logic [34]. This translation is achieved as follows: $f_{P\rightarrow\Omega}(\pi):S\times A\rightarrow\Omega~|~P.$ As explained in Sec. II-C, probability $p$ translates to opinion $\omega:p\mapsto(p,1-p,0,p).$
@@ V-C2 Policy shaping by opinion fusion
To fuse opinions, [34] defines several fusion operators that are classified based on the situation. In this work, the agent and advisor formulate their opinions separately, and it is fair to assume they will not compromise. Thus, we use the Belief Constraint Fusion (BCF) operator, as it is most appropriate when compromise is not sought, and those giving opinions have made up their minds. Joint opinion ${\omega}^{\odot}=({b}^{\odot},{d}^{\odot},{u}^{\odot},{a}^{\odot})$ under BCF is calculated by first finding the harmony and conflict of opinions: $\textit{Harmony}=b_{1}u_{2}+b_{2}u_{1}+b_{1}b_{2}$; $\textit{Conflict}=b_{1}d_{2}+b_{2}d_{1}$. Then, the components of the joint opinion are calculated as follows.
@@ V-D Transforming opinion-infused policies from the certainty domain to the probability domain
The agent’s policy now reflects the advice, and the agent may begin the exploration process.